OpenStudy (anonymous):

A sample of gas occupies a volume of 445 mL at a temperature of 274 K. At what temperature will this sample of gas occupy a volume of 499 mL? Show all of the work used to solve this problem.

OpenStudy (anonymous):

@Somy Can you help? :3

OpenStudy (somy):

charles's law formula is V1/T1 = V2/T2

OpenStudy (anonymous):

I need the formula thanks! I'm going to do the work and can you double check it for me?

OpenStudy (somy):

sure :))

OpenStudy (anonymous):

\[\frac{ 445 }{ 274 } * \frac{ 499 }{ t }\] \[445(t) = 499 *274\] 445(t) = 136726 \[\frac{ 445(t) }{ 445} = \frac{ 136726 }{ 445}\] t = 307 k

OpenStudy (somy):

yeah thats right :)

OpenStudy (anonymous):

Omg, yay! Can you help me set up another one?

OpenStudy (somy):

yup yup :) but open a new question :)

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

This is a dishonest way to complete you exam...